Two electronic components of a missile system work in harmony for the success of the total system. Let X and Y denote the life in hours of the two components. The joint density of X and Y is −y(1+x) ye , x, y ≥ 0, f(x, y) = 0, elsewhere.

(a) Give the marginal density functions for both random variables.

(b) What is the probability that the lives of both components will exceed 2 hours?
Answer #1

Solution :- Giveu, f(x, y) = f de o x, 47,0 lo else a) Marginal distribution of 8(x) = f (x, G) dy - de y Cita) ay y €1+x)=zMarginal distribution of y o nco) - S & (mix) dx = open at ofecto das ye Seat sh(y) = { e o o b) P(*7 , y>2) go, ; else geoli